the sum of two integers, a and b (a>b). the smaller minus the larger is -746. find the value of the integers a and b.
Answer by checkley71(8403)

IF YOU'VE DEFINED THE PROBLEM CORRECTLY. THERE ARE MANY SOLUTIONS THUS:
A>B THEN B IS THE SMALLER NUMBER &
B-A=-746 THEN IF B=1 THEN
1-A=-746
-A=-746-1
A=747
SETTING B=242 WE GET
242-A=-746
-A=-746-242
A=988